A Radio Tuning Circuit

A radio can be tuned into a particular station frequency byadjusting the capacitance in an L-C circuit. Suppose thatthe minimum capacitance of a variablecapacitor in a radio is4.14 pF.a) What is the inductance Lof a coil connected to this capacitor ifthe oscillation frequencyof the L-C circuit is 1.70 MHz, corresponding to one end of the AM radio broadcast band,when the capacitor is set to itsminimum capacitance?b) The frequency at the other end of the broadcast band is0.538 MHz. What is the maximum capacitance C_max of the capacitor if the oscillation frequency isadjustable over the range of the broadcastband?
Answer #1

Hi,
Given data:
The minimum capacitance of a variable capacitor in a radio = 4.14pF =4.14x10-12F.
( a ) Frequency of the L-C circuit = 1.70MHz = 1.7x106
Findinductance L ?

In this case, if we want to tune any frequency ( i.eresonant frequency ) conditionis,
Xc =XL
∴ 1 / 2πfC =2πfL
1 / ( 2 x3.14 x 1.7 x 106 x 4.14 x 10-12 ) = 2 x3.14 x 1.7 x 106 x L

Simplifyand get L value = 1.48 mH


( b ) The other end of the broadcast band is0.538 MHz
FindCmax = ?
L = 1.48mH ( Taken from above problem (Fixed))

Xc = XL ; 1 /2πfC = 2πfL

C= 1 / ( 4 π2f2 L )
Cmax = 1 / ( 4π2 x (0.538 x 106 )2x 1.48 x 10-3 )
Cmax =0.237 x 10-9 = 0.237 nF

I hope this will helps you.

answered by: ramla A